2. The cash budget: Sit Down Corporation has a cash balance of $32,500 on April 1. The company must maintain a minimum cash balance of $30,000. During April, expected cash receipts are $48,500. Cash disbursements during the month are expected to total $56,100. Ignoring interest payments, how much will the company need to borrow during April?
Answer #1

Answer:

Amount($)
A Beginning Cash Balance - Apr 1                  32,500
B Cash receipts                  48,500
C = A + B Total Cash Available                  81,000
D Cash Disbursements                  56,100
E = C - D Excess of Cash available over disbursements                  24,900
F Ending balance desired                  30,000
G = F - E Amount to be borrowed                    5,100